# Network Formation and Dynamics Among Multi-LLMs

### v3

@@ Abstract.
Social networks shape opinions, behaviors, and information dissemination in human societies. As large language models (LLMs) increasingly integrate into social and professional environments, understanding their behavior within the context of social interactions and networks becomes essential. Our study analyzes LLMs’ network formation behavior to examine whether the dynamics of multiple LLMs are similar to or different from human social dynamics. We observe that LLMs exhibit key social network principles, including preferential attachment, triadic closure, homophily, community structure, and the small-world phenomenon, when asked about their preferences in network formation. We also investigate LLMs’ decision-making based on real-world networks, revealing that triadic closure and homophily have a stronger influence than preferential attachment and that LLMs perform well in network formation predictions. Overall, our study opens up new possibilities for using LLMs in network science research and helps develop socially aware LLMs by shedding light on their social interaction behaviors and exploring their impacts on social dynamics.
@@ Introduction
Recent progress in large language models (LLMs), such as GPT (OpenAI, 2023) and Llama 2 (Touvron et al., 2023), have shown promising developments in AI techniques and their integration into real-life applications. It is thus crucial to comprehend AI actions to ensure they align with human expectations, mitigate potential risks, and maximize their benefits. Misaligned AI actions may lead to unintended consequences, such as biased decision-making, fairness issues, and the miscoordinative or non-cooperative behavior (Rahwan et al., 2019). Recently, researchers have started to apply social science methodologies, such as methods analogous to laboratory experiments (Horton, 2023; Aher et al., 2023; Veselovsky et al., 2023; Manning et al., 2024), agent-based modeling (Park et al., 2018; Gao et al., 2023; He et al., 2023; De Marzo et al., 2023; Fatemi et al., 2023; Perozzi et al., 2024), and qualitative methods, to study LLMs. These methods not only reveal the capabilities and interpretability of LLMs but also suggest their potential for applications in social science (Horton, 2023; Park et al., 2023; Chen et al., 2023; Leng et al., 2023).
In human societies, social networks play a crucial role in shaping individual behaviors, preferences, and connections, as well as influencing the diffusion of information and norms across communities (Rogers, 2003; Bakshy et al., 2012; Fowler and Christakis, 2008; Banerjee et al., 2013; Yuan et al., 2018). LLMs have shown great potential in social contexts, notably as intelligent personal assistants that facilitate social and prosocial interactions (see, e.g., (Papachristou et al., 2023; Chopra and Haaland, 2023; Veselovsky et al., 2023)). However, less is known about how LLMs’ innate behaviors and preferences align with human network formation principles (Jackson, 2008; Zhou et al., 2023; Park et al., 2023). This is particularly crucial, as it sheds light on the potential of these models to shape and be shaped by the networks of human relationships, which is a fundamental aspect of social systems.
@@
In simulations conducted using synthetic networks, we have the following findings: First, we observe that LLMs can generate networks exhibiting preferential attachment and following scale-free degree distributions. Moreover, LLMs demonstrate strong tendencies toward triadic closure and exhibit tendency of homophily. These tendencies further explain the emergence of community structures in networks formed by multi-LLMs. Additionally, we show that LLM agents exhibit dynamics consistent with small-world phenomena akin to those observed in the Watts-Strogatz model (Watts and Strogatz, 1998).
Next, we compare between these principles to identify the key factors driving network dynamics in real-world network formation contexts. To achieve this, we consider real-world networks and ask each LLM to simulate a node in the network. We provide information on their network structure and ask for their decisions for the next link formation or dissolution step. We approach this problem as a discrete choice modeling problem (Liben-Nowell and Kleinberg, 2003; Overgoor et al., 2019). In this real-world setting, we find that LLMs’ choices are primarily weighted on homophily and triadic closure, compared to preferential attachment, and that LLMs substantially exceed random guessing when predicting new links. Additionally, the resulting choices of LLM agents strengthen the network’s community structure.
Our study contributes to the literature on LLMs in multiple ways. First, it advances the discussion on aligning LLMs with human preferences, which has potential applications in developing socially aware LLMs for real-life interactions. Insights into LLMs’ innate behaviors in network formation can inform the design of LLMs for effective collaboration in social and professional settings. Second, our study could contribute to understanding the impact of LLMs on social dynamics and norms, such as their role in shaping online communities’ interaction patterns or influencing the adoption of new communication norms in professional settings. Finally, our work has direct applications to synthetic dataset generation, a problem important for the network science community. By controlling for a few parameters, such as the temperature, our work can be utilized to generate a variety of datasets, as well as generate synthetic datasets from real-world data to protect privacy.
(a) Structure-based
(b) Degree-based
Figure 1. Results for Principle 1 (preferential attachment): We display simulated networks comprising 200 nodes across different temperatures. For structure-based simulations (Figure 1(a)), the multi-LLM setup was given neighborhood information $\{N_{j,t}:j\in V_{t}\}$, while for degree-based simulations (Figure 1(b)), node degree data $\{d_{j,t}:j\in V_{t}\}$ was provided ($V_{t}$ corresponds to the vertex set of the network $G_{t}$ at round $t$). The degree distributions were then compared with those of a BA network with $n=200$ nodes. (a): Power-law degree distributions are evident ($P>0.5$, K-S test), with the networks at a temperature of 1.5 closely resembling the BA model ($P>0.1$, K-S test). (b): With degree information only, the networks form more unrealistic star-like structures, diverging from scale-free configurations and more closely mirroring a core-periphery network structure.
@@ Results
In our study, we leveraged state-of-the-art LLMs through the OpenAI Application Programming Interface (API). Our methodology involved constructing multiple LLMs, each acting as an independent agent within separate conversational threads. Each LLM agent was prompted with network information about its peers and faced with decision-making scenarios. For synthetic network simulations, we employed GPT-3.5 agents, whereas GPT-4 agents were used for real-world network analysis. [^1] We utilized the gpt-3.5-turbo model for GPT-3.5 and the gpt-4-1106-preview model for GPT-4, both provided by OpenAI. Notably, both GPT-3.5 and GPT-4 yielded consistent results.
We adopt a multi-LLM framework, emphasizing the multi-agent aspect of our experimental design, which shares similar spirits with recent research in the field (Leng et al., 2023; Akata et al., 2023; Brookins and DeBacker, 2023; Aher et al., 2023; De Marzo et al., 2023). To assess the robustness of our findings, we varied the ‘temperature’ settings in LLMs, which modulate the output variability, in a dynamic network environment. Additionally, we adopted an interview-like method to probe the rationale behind the LLMs’ decision-making processes, enriching our understanding of their behavior in network simulations.
In our analysis, we begin with an examination of three fundamental micro-level network principles: preferential attachment, triadic closure, and homophily. Subsequently, we shift our focus to two macro-level network phenomena: community structure and the small-world phenomenon. Finally, we assess the resilience of these principles as LLM agents interact with real-world networks. This allows us to compare and contrast the network formation preferences between LLMs and humans.
@@
In our study, we investigate whether LLMs manifest the preferential attachment mechanism. Starting with an empty network, we introduce nodes sequentially at each time step $t$. A new node $i_{t}$ is added during time step $t$, with the network’s existing nodes providing context about their connections. We generated networks with $n=200$ nodes, a size large enough to observe meaningful patterns in the degree distribution. Our simulations include two distinct types: structure-based, where the network’s structure was provided, and degree-based, where only the degree information of potential connections was given. [^2] For details on how features were represented in the prompts, refer to SI Section 5.
The graphical outputs from our structure-based simulation, as illustrated in Figure 1(a), reveal a distinct pattern where a small number of nodes possess a high number of connections, whereas the majority exhibit fewer links. This pattern suggests a scale-free degree distribution, characterized by the probability $\pi(d)$ that a node has degree $d$, mathematically represented as:
(1) $\pi(d)\propto d^{-\gamma},\quad\text{where }\gamma>1.$
Interestingly, we also observe that the number of highly connected “super nodes” increases with temperature.
[…]
Furthermore, in degree-based simulations in a similar spirit of ((De Marzo et al., 2023); see Figure 1(b)), we noted formations resembling a “star” around central nodes. These distributions deviate from the scale-free patterns seen in statistical models like the BA model ($P<0.001$, two-sample K-S test), aligning more closely with core-periphery structures (Borgatti and Everett, 2000).
To interpret the LLM agents’ decision-making, we analyzed the reasoning they provided, focusing on the structure-based results (Figure 1(a)) [^4] The degree-based results’ reasoning is not included as decisions were based solely on degree information.. Two primary reasons emerged, varying with temperature (detailed in SI Section 3):
  • Well-Connectedness: Relating to the node’s degree (number of connections).
  • Mutual Friends: Relating to the number of shared connections.
Interestingly, as the temperature increased, the prevalence of the well-connectedness rationale decreased, while the mutual friends’ rationale increased.
In conclusion, our analysis demonstrates that LLMs, when provided with network structure, can replicate the preferential attachment phenomenon.
@@
(a) Structure-Based
(b) Common-Neighbor-Based
Figure 2. Results for Principle 2 (triadic closure). (a) Figure 2(a) shows the resulting networks created by triadic closure when the intersection of the neighborhoods of the query node and each alternative is provided. The node colors correspond to the stochastic block model (SBM) groups to which each node belongs. The red edges correspond to the newly created edges between nodes of the red cluster, the blue edges correspond to the newly created edges between nodes of the blue cluster, and the orange edges correspond to edges created between the two clusters. The newly formed edges are drawn with a larger line width. We observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test). (b) Figure 2(b) shows the same networks with the only change that instead of the intersection of neighborhoods between the query node and each alternative, we provide the number of common neighbors (i.e., the size of the intersection) between the query node and each alternative. Similarly, we observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test). The error bars correspond to 95% confidence intervals.
@@
To investigate triadic closure, we employ an assortative stochastic block model (SBM) (Newman, 2003) to create an initial network $G_{1}$ with $n$ nodes divided into two equal-sized clusters $A$ and $B$. Connections within each cluster are formed with a probability of $0.5$, while inter-cluster connections occur with a probability of $0.1$. This setup mirrors our assumption that nodes within the same cluster are more inclined to connect due to a higher number of shared neighbors. In subsequent time steps, we then examine each node $i$, considering the intersection of neighborhoods of $i$’s non-neighbors or the number of mutual friends between $i$ and any non-neighbor $j$ [^5] It’s worth noting that similar outcomes arise when providing neighbors instead of common neighbors.
We conducted ten simulations with $n=50$ nodes to facilitate clear visualization and ensure statistical significance [^6] Choosing $n=50$ instead of a larger number like $n=200$ aids in visualization and maintains statistical significance.. To contrast LLM-based network formation, we refer to a random null model, selecting nodes at random for connection.
For evaluating triadic closure, we utilize two metrics: marginal transitivity and probability of edge formation within the same community. Marginal transitivity ($D$) represents the change in the ratio of closed triangles to all triads, transitioning from the initial network $G_{T}$ to the SBM-generated $G_{1}$:
@@
where a large positive $D$ indicates a strong triadic closure tendency.
The probability of forming an edge within the same community ($\hat{p}$) is calculated by the ratio of edges in $G_{T}\setminus G_{1}$ (newly formed edges) connecting nodes within the same cluster:
@@
In Figure 2(a), sample networks are displayed, with the upper panel showing networks where the entire structure is provided and the lower panel showing those with only common neighbor numbers provided. Nodes are color-coded to indicate their cluster memberships in the SBM, with red and blue edges within clusters and orange edges between clusters. Newly formed edges are highlighted with thicker lines.
Marginal transitivity, presented in the fourth panel of each row, shows a statistically significant increase across all temperatures ($P<0.001$, t-test), and compared to the Random model, the increase is notable ($P<0.001$, t-test).
Regarding the probability of within-community network formation (last panels in each row), we found that across all temperatures, this probability significantly exceeds that in the random network formation scenario ($P<0.001$, t-test against Random), indicating an LLM tendency towards triadic closure.
These findings affirm that LLM agents tend to follow triadic closure. We investigate the reasons provided by LLMs for their decisions in SI Section 3.777Reasons for decisions based on the number of common neighbors are not provided, as these decisions are solely based on this information. The analysis revealed that many decisions were influenced by the number of mutual friends.
@@
To test whether LLM agents exhibit homophily, we perform the following experiment: We generate nodes with randomly generated attributes regarding a hobby (randomly chosen among three hobbies), a favorite color (randomly chosen among three colors), and a location within the US (randomly chosen among three US locations) and provide the attributes of the other nodes and the node’s own attributes, and each node is tasked to form up to $\delta=5$ links with others. For each node $i$, we provide it with the features $x_{j}$ of all non-neighbors $j$ of $i$. The seed network is taken to be the empty graph.
We run ten simulations for networks with $n=50$ nodes and $\delta=5$. We compare with the Random null model, which chooses $\delta=5$ links at random.
To measure the presence of homophily, we measure the attribute assortativity coefficient of each of the three features (hobby, color, location). For each property $P$ which takes $K$ distinct values $P_{1},\dots,P_{K}$ (indexed by $k$ or $l$), its assortativity coefficient $R$ is defined as
@@
In Figure 3(a), we observe a consistent pattern of positive assortativity across all temperature ranges, which is higher for lower temperatures. Moreover, the observed assortativity at all temperatures significantly differs from Random, which has assortativity close to 0 ($P<0.001$, t-test).
Moreover, we present the probability of each reason being mentioned in SI Section 3. We find that shared hobbies appear to be much more important than shared location or favorite color. This is consistent with our intuition that, similarly to humans, LLMs care more about hobbies than other superficial attributes when forming social connections. In token, we conclude homophily emerges for the networks generated by LLMs.
@@
(a) Assortativity Coefficients
(b) Network Instances
(c) Communities Statistics
Figure 3. Results for Principle 3 (homophily). (a) Figure 3(a) shows the assortativity coefficients for each network attribute created based on Principle 3. Each node can form at most $\delta=5$ links with the other nodes based on their attributes. We observe that for all temperatures, we have positive assortativity ($P<0.001$, t-test with 0), which indicates homophily. The assortativities are significantly different for all temperatures from random or heterophilous ($P<0.001$, t-test comparing with Random). (b) Networks created based on Principle 3. Each node can form at most $\delta=5$ links with the other nodes based on their attributes. The edge widths correspond to the number of common attributes shared between the nodes. The node colors correspond to the communities found by running the Louvain algorithm. (c) Louvain modularity, number of communities, and community sizes of networks created based on Principle 3. Lower temperatures yield higher modularity ($P<0.001$, t-test) and fewer and larger communities ($P<0.001$, t-test). For smaller temperatures, the communities are more “well-separated”. The error bars correspond to 95% confidence intervals.
@@
The community structure of networks refers to the organization of nodes or individuals within a network into distinct and densely interconnected groups or clusters (Newman and Girvan, 2004; Newman, 2006; Blondel et al., 2008; Clauset et al., 2004). Identifying community structures is crucial for understanding the overall dynamics of a network, as it reveals patterns of relationships and interactions that might not be apparent at the global level.
@@
In Figure 3(b), various colors represent the communities identified by the Louvain algorithm at different temperatures. Notably, communities appear more distinct at lower temperatures, likely due to reduced randomness in decision-making at these temperatures. Additionally, Figure 3(c) presents the distribution of Louvain modularity values across simulations, indicating consistent community structure with positive modularity at all temperatures, confirmed by a t-test against a modularity of $Q=0$ for a random graph ($P<0.001$). Moreover, lower temperatures lead to fewer, larger communities as per the Louvain algorithm’s findings ($P<0.001$, t-test).
@@
Figure 5. Fitted results for Principle 5 (small world). Top: Regression plots relating average shortest path length ($L$) and average clustering coefficient ($C$) with $n$ for $\beta\in\{0.25,0.5,0.75\}$ and $k=5$. Value $a$ in legends represents the slope of the regression lines. Bottom: Estimated values $\hat{\beta}$ of $\beta\in\{0.25,0.5,0.75\}$ for LLM-generated networks based on matching the average clustering coefficient and difference in the average shortest path between LLM-generated networks and Watts-Strogatz with the estimated rewiring probability $\hat{\beta}$. We report the $P$-values of the t-test comparing the average shortest path length of the LLM-generated networks and the average shortest path length of the Watts-Strogatz graphs with rewiring probability $\hat{\beta}$. *: $P<0.05$; **: $P<0.01$, and ***: $P<0.001$.

### v4

@@ Abstract.
Social networks fundamentally shape human opinions, behaviors, and the dissemination of information. As large language models (LLMs) like GPT, Claude, and Llama increasingly integrate into social and professional settings, understanding their behavior in the context of social interactions and network formation becomes essential. This study develops a framework to systematically examine whether the network formation behaviors of multiple LLMs approximate certain aspects of human network dynamics. By simulating interactions among LLM agents across various model families, we observe that these models consistently exhibit key patterns associated with social network principles—including preferential attachment, triadic closure, homophily, community structure, and the small-world phenomenon—when forming networks. Moreover, LLMs adapt their network formation strategies based on each network’s characteristics, reflecting the context-dependent nature of human behavior: in Facebook networks, they prioritize triadic closure and homophily, mirroring close-knit friendships; in phone networks, homophily and preferential attachment dominate, capturing personal and professional connections, while in employment networks, LLMs favor heterophily and high-degree connections, aligning with career advancement dynamics. These results open new avenues for using LLMs in network science research, with potential applications in agent-based modeling and synthetic network generation.
@@ Introduction
Recent progress in large language models (LLMs), such as GPT (OpenAI, 2023), Claude (Anthropic, 2024), and Llama (Touvron et al., 2023), has shown promising developments in AI techniques and their integration into real-life applications. It is thus crucial to comprehend AI actions to ensure they align with human expectations, mitigate potential risks, and maximize their benefits. Misaligned AI actions may lead to unintended consequences, such as biased decision-making, fairness issues, and the miscoordinative or non-cooperative behavior (Rahwan et al., 2019). Recently, researchers have started to apply social science methodologies, such as methods analogous to laboratory experiments (Horton, 2023; Aher et al., 2023; Veselovsky et al., 2023; Manning et al., 2024), agent-based modeling (Park et al., 2018; Gao et al., 2023; He et al., 2023; De Marzo et al., 2023; Fatemi et al., 2023; Perozzi et al., 2024), and qualitative methods (Chew et al., 2023), to study LLMs. These methods not only reveal the capabilities and interpretability of LLMs but also suggest their potential for applications in social science (Horton, 2023; Park et al., 2023; Chen et al., 2023; Leng et al., 2023).
In human societies, social networks play a crucial role in shaping individual behaviors, preferences, and connections, as well as influencing the diffusion of information and norms across communities (Rogers, 2003; Bakshy et al., 2012; Fowler and Christakis, 2008; Banerjee et al., 2013; Yuan et al., 2018). LLMs have shown great potential in social contexts, notably as intelligent personal assistants that facilitate social and prosocial interactions (see, e.g., (Papachristou et al., 2023; Chopra and Haaland, 2023; Veselovsky et al., 2023)). However, less is known about how LLMs’ behaviors and preferences align with human network formation principles (Jackson, 2008; Zhou et al., 2023; Park et al., 2023). This is particularly crucial, as it sheds light on the potential of these models to shape and be shaped by the networks of human relationships, which is a fundamental aspect of social systems.
@@
In synthetic network simulations, LLMs displayed preferential attachment, homophily, and triadic closure, resulting in the formation of community structures and small-world dynamics. More notably, in real-world social network simulations, we find that LLMs prioritized triadic closure and homophily over preferential attachment when forming new links, indicating a strong preference for connecting with similar nodes or shared acquaintances. Additionally, in a telecommunication network, LLMs tended to prioritize homophily and preferential attachment over triadic closure, and in a company network, the agents who corresponded to employees formed links frequently with managers, which showcases behavior that is consistent with human social mobility principles.
Generally, LLMs not only exhibit fundamental social network formation principles in synthetic simulations but also adapt their strategies based on the context of real-world networks, mirroring human social behaviors specific to each setting.
As LLM technology continues to evolve, our study serves as an early exploration of their potential in social network studies, with several significant implications for future studies. First, our study demonstrates the potential of LLMs for agent-based modeling. By simulating decision-making processes that approximate human-like behavior across various network settings, LLMs can provide valuable insights into the emergence of social phenomena. Although these models are still in early stages of development, they offer an intriguing framework for studying and designing systems that can mimic key aspects of real-world dynamics. This opens up possibilities for applying LLMs to explore and understand complex behaviors in social, professional, and collaborative environments. Second, our work highlights the potential of LLMs for synthetic dataset generation, a critical area in network science. Although the accuracy of LLM-based predictions is not yet perfect like all other link prediction models, this approach is particularly valuable in scenarios where privacy concerns limit access to real-world data. By simulating realistic datasets that capture important network properties, LLMs can facilitate research and experimentation without compromising sensitive information.
@@ Results
In this study, we investigated whether LLMs exhibit fundamental principles of network formation observed in human social networks. By simulating multiple LLM agents acting independently within separate conversational threads, we examined their behaviors in decision-making scenarios involving network connections. We focused on three micro-level network principles – preferential attachment, triadic closure, and homophily – and two macro-level phenomena—community structure and the small-world effect. To assess the robustness of our findings, we varied the temperature settings of different LLM models, including GPT-3.5-turbo, GPT-4o Mini, Llama 3 (70b-instruct), and Claude 3.5 Sonnet. We also experimented with different environmental prompts (e.g., friendship, collaboration, community) to test prompt sensitivity. Additionally, we employed an interview-like method to probe the LLMs’ decision-making rationale and conducted experiments using Chain-of-Thought (CoT) reasoning (Wei et al., 2022) (the experiments are deferred to Appendix C). Finally, we extended our analysis to real-world networks, including a social media friendship network, a telecommunication network, and a company collaboration network, to compare the network formation preferences between LLMs and humans.
More information about the experimental procedure can be found in the Methods and Materials section of this paper.
@@
(a) Probability of connecting to top-$k$ nodes for different models, temperatures, and environments
(b) Power law fits ($\hat{\gamma}$) and standard errors for different models, temperatures, and environments
(c) Network instances generated by GPT-3.5 agents
Figure 1. Results for Principle 1 (preferential attachment) The multi-LLM setup was given neighborhood information $\{N_{j,t}:j\in V_{t}\}$. Top: Probability of connecting to top-$k$-degree nodes for varying model (temperature is fixed to 1.0 and environment to baseline), temperature (model fixed to GPT-3.5 and environment to baseline) and environment (model fixed to GPT-3.5 and environment temperature to 1.5) for networks generated according to Principle 1 with $n=200$ nodes. Middle: Power Law exponents and standard errors for varying model, temperature, and environment. Bottom: Simulated networks. Power-law degree distributions are evident ($P>0.5$, K-S test), with the networks at a temperature of 1.5 closely resembling the Barabási-Albert model ($P>0.1$, K-S test) for GPT-3.5 agents.
@@
To test if LLM agents exhibit preferential attachment, we simulated network growth by sequentially adding nodes to an initially empty network. Each new node was prompted with information about existing nodes, and the person to connect with was decided. We generated networks with $n=200$ nodes to observe meaningful degree distributions [^1] Note that we provide the full network structure in the prompt, so models are not inherently biased toward forming links with the highest-degree nodes..
On a micro-scale, Figure 1(a) illustrates the probability of connecting to a top-$k$ node as a function of its degree percentile ($k/n$). To demonstrate the tendency toward preferential attachment, we compare these probabilities to a null model assuming random connections (represented by dashed lines), where the likelihood of connecting to a top-$k$ node is simply $k/n$. Our findings reveal that all models prefer connecting to higher-degree nodes. Notably, GPT-3.5 exhibits a weaker preference, while other, arguably more capable models, show an even stronger inclination toward preferential attachment. Using GPT-3.5 as an example, we examine the effect of temperature – a parameter controlling the variability of model output – on this tendency. At lower temperatures, the model makes fewer stochastic choices and, as a result, is more likely to connect to high-degree nodes. We also vary the prompt to explore the influence of “environment”-contextual settings such as school, work, or community. The results show slight variations compared to the baseline (GPT-3.5 with temperature = 1.5), yet the tendency for preferential attachment persists across environments. In all cases, the observed curves lie above the null model, underscoring the presence of preferential attachment.
Next, we investigate the degree distribution of the resulting graphs. As shown in Figure 1(c), the resulting networks display a pattern where a few nodes have many connections while most have few, indicative of a scale-free distribution, with form:
(1) $\pi(d)\propto d^{-\gamma},\quad\text{where}\quad\gamma>1.$
We estimated the exponent $\gamma$ for different models and temperatures. Our analysis reveals several notable patterns in the networks generated by LLM agents under different conditions. First, models newer than GPT-3.5 exhibit a slightly larger $\hat{\gamma}$ than GPT-3.5. This implies that these models display a stronger tendency toward preferential attachment and the formation of hubs. Second, as the temperature increases the power-law exponent $\hat{\gamma}$ generally becomes larger. This indicates that higher temperatures introduce more variance in node connectivity, leading to degree distributions with heavier tails. Third, the environmental context significantly affects the value of $\hat{\gamma}$. For example, when the network is framed within a “school” environment, the exponent increases, suggesting a more uniform distribution of connections and fewer highly central nodes.
Finally, while the prompts we have utilized thus far have provided the model with the complete existing network structure, we also explore an alternative scenario: what happens if agents are supplied solely with the degree of other alternatives, without access to the network’s full structure? As detailed in Appendix B, our findings reveal that limiting agents to degree information alone also leads to notable structural differences in the networks that emerge (cf Figure S.9). Thus, degree information alone yields more restrictive structures than providing the agents with the full topological information (i.e. the neighbors).
The findings highlight the practical potential of LLMs in modeling complex networks, such as social, economic, or biological systems, by leveraging their ability to simulate preferential attachment and scale-free distributions. These models can be used to study real-world phenomena like information diffusion, hub formation, or connectivity patterns under varying conditions. Additionally, the sensitivity of network structures to parameters like temperature and context underscores the importance of prompt design in steering outcomes, making LLMs versatile tools for tailored simulations.
@@
(a) Probability of connecting to top-$k$ for different models, temperatures, and environments
(b) Marginal transitivity ($D$) and probability of an edge within a community ($\hat{p}$) for different models, temperatures, and environments
(c) Network instances by GPT-3.5 agents
Figure 2. Results for Principle 2 (triadic closure). Top: Probability of connecting to top-$k$ nodes (in terms of common neighbors) for varying model (temperature is fixed to 1.0 and environment to baseline), temperature (model fixed to GPT-4 Mini and environment to baseline) and environment (model fixed to GPT-4 Mini and environment temperature to 0.5) for networks generated according to Principle 2 ($n=50$, 10 simulations for each model, environment and temperature). Middle: Marginal transitivity ($D$) and probability of an edge within a community ($\hat{p}$) for networks generated according to Principle 2 in different models, temperatures, and environments. Bottom: The figure shows the resulting networks created by GPT-4 Mini, according to Principle 2 when the intersection of the neighborhoods of the query node and each alternative is provided. The node colors correspond to the groups to which each node belongs. The bold edges (red or blue) correspond to the newly created inter-cluster edges, and the orange edges correspond to the new intra-cluster edges.
@@
To investigate triadic closure, we employ an assortative stochastic block model (SBM) (Newman, 2003) to create an initial network $G_{1}$ with $n$ nodes divided into two equal-sized clusters $A$ and $B$. Connections within each cluster are formed with a probability of $0.5$, while inter-cluster connections occur with a probability of $0.1$. This setup mirrors our assumption that nodes within the same cluster are more inclined to connect due to a higher number of shared neighbors. In subsequent time steps, we then examine each node $i$, considering the intersection of neighborhoods of $i$’s non-neighbors [^2] Similar outcomes arise when providing neighbors instead of common neighbors..
We conducted ten simulations with $n=50$ nodes to facilitate clear visualization and ensure statistical significance [^3] Choosing $n=50$ instead of a larger number like $n=200$ aids in visualization and maintains statistical significance..
On a micro-scale, Figure 2(a) illustrates the probability of connecting to a top-$k$-percentile node as a function of the number of common neighbors. The dashed lines represent the results of null models, where connections are chosen randomly; which corresponds to the probability of connecting to a top-$k$ percentile node in terms of the common neighbors being $k/n$. Our findings reveal that, across all models, there is a consistently higher probability of forming links with nodes that share more common neighbors. Unlike the behavior observed in preferential attachment, temperature does not appear to severely impact this probability. This tendency to form links with nodes that have more common neighbors is consistent across various contexts, including school, work, and community environments. These results suggest that the triadic closure tendency is a robust phenomenon, persisting across different model families, configurations, and environments.
Then, for evaluating triadic closure on the network (macroscopic) level, we utilize two metrics: marginal transitivity and probability of edge formation within the same community. Marginal transitivity ($D$) represents the change in the ratio of closed triangles to all triads, transitioning from the initial network $G_{T}$ to the SBM-generated $G_{1}$:
@@
where a large positive $D$ indicates a strong triadic closure tendency. The probability of forming an edge within the same community ($\hat{p}$) is calculated by the ratio of edges in $G_{T}\setminus G_{1}$ (newly formed edges) connecting nodes within the same cluster:
@@
Marginal transitivity ($D$), presented in Figure 2(b), demonstrates a statistically significant increase across all models, temperatures, and environments, underscoring the robust nature of triadic closure. Similarly, the probability of within-community edge formation ($\hat{p}$), shown in Figure 2(b), consistently exceeds $0.5$, reaffirming the tendency of nodes to form new connections within their respective clusters.
In Figure 2(c), sample networks from GPT-3.5 are displayed, with the upper panel showing networks where the entire structure is provided and the lower panel showing those with only common neighbor numbers provided. Nodes are color-coded to indicate their cluster memberships in the SBM, with red and blue edges within clusters and orange edges between clusters. Newly formed edges are highlighted with thicker lines.
In summary, these findings show that most LLMs exhibit a consistent tendency for triadic closure across various configurations, temperatures, and environments. This behavior mirrors human network dynamics, highlighting the models’ ability to simulate realistic social and structural networks and reinforcing their alignment with social principles observed in real-world communities.
@@
(a) Assortativity and Louvain Modularity with different LLM models and environments
(b) Network instances generated by GPT-3.5 agents
Figure 3. Results for Principle 3 (Homophily) and Principle 4 (Community structure due to homophily). Top: Assortativities and Louvain modularity according to Principle 3 ($n=50$, 5 simulations for each row) in different environments (school, work, community) using different models. The statistical significance is $P<0.001$ for all t-tests (comparing with 0). Bottom: Network instances and community structure.
Figure 4. Effect of distractor features (favorite color and lucky number) on homophily.
@@
To test whether LLM agents exhibit homophily, we perform the following experiment: We generate nodes with randomly generated attributes regarding a hobby (randomly chosen among three hobbies), a favorite color (randomly chosen among three colors), and a location within the US (randomly chosen among three US locations) and provide the attributes of the other nodes and the node’s own attributes, and each node is tasked to form up to $\delta=5$ links with others. For each node $i$, we provide it with the features $x_{j}$ of all non-neighbors $j$ of $i$. The seed network is taken to be the empty graph. We run ten simulations for networks with $n=50$ nodes and $\delta=5$.
To evaluate homophily, we calculate the attribute assortativity coefficient for each of the features. For each property $P$ which takes $K$ distinct values $P_{1},\dots,P_{K}$ (indexed by $k$ or $l$), its assortativity coefficient $R$ is defined as
@@
From Figure 3(a), we observe that different attributes exhibit varying levels of assortativity. First, homophily is present across all LLMs—regardless of the specific model or configuration (e.g., temperature settings), all show positive assortativity for all four attributes. This aligns with human societies, where homophily is a primary driver of network formation (McPherson et al., 2001) [^4] As an additional robustness check, we also tested mutual agreement connections. In that setting, after a node $j$ is chosen by node $i$, $j$ has to confirm the creation of the link from itself to $i$ ($j\to i$). We ran several experiments with different models and temperatures and we found the results not to be affected, namely, the proposed connections were always bilateral..
Moreover, to test the effect of the features on homophily as indicated by the assortativity coefficient for each attribute, we introduce a distractor feature, which corresponds to a lucky number that is randomly chosen between 0 and 9. We repeat the simulations for all models and measure the effect of each feature on Figure 4. We show that lucky numbers consistently show lower assortativity coefficients, indicating they are less considered when forming homophilous connections. This is consistent with our prior expectation that humans typically do not prioritize shared lucky numbers when establishing relationships.
Surprisingly, even though the lucky number does not seem to impact homophily much, the favorite color exhibits a similar level of homophily as hobbies. One might expect that hobbies, being substantive interests, would have a stronger influence on social connections than favorite colors, which are more arbitrary preferences. However, this finding aligns with the social identity theory and the minimal group paradigm (Tajfel et al., 1971). According to this paradigm, even minimal and arbitrary group distinctions – such as a preference for certain colors – can lead to in-group favoritism and influence social connections. This suggests that LLM agents, akin to humans, may form connections based on even trivial shared attributes, reflecting inherent tendencies toward group formation based on minimal commonalities.
All in all, LLMs can capture and reproduce subtle human social behaviors, not just linguistic patterns. This underscores their potential as powerful tools for social simulation. However, these findings may also raise important considerations regarding bias, fairness, and the ethical design of AI systems (cf. Discussion Section).
@@
The community structure of networks refers to the organization of nodes or individuals within a network into distinct and densely interconnected groups or clusters (Newman and Girvan, 2004; Newman, 2006; Blondel et al., 2008; Clauset et al., 2004). Identifying community structures is crucial for understanding the overall dynamics of a network, as it reveals patterns of relationships and interactions that might not be apparent at the global level. [^5] As an example, we present only the results from GPT-3.5 for Principle 4 (Community Structure) and Principle 5 (Small-World).
@@
In Figure 3(b), various colors represent the communities identified by the Louvain algorithm at different temperatures for GPT-3.5. Notably, communities appear more distinct at lower temperatures, likely due to reduced randomness in decision-making at these temperatures.
Figure 3(a) presents the distribution of Louvain modularity values across simulations accross different LLM models and different environments, indicating consistent community structure with positive modularity at all temperatures, confirmed by a t-test against a modularity of $Q=0$ for a random graph ($P<0.001$).
@@
(g) Regression plot for different models and environments for $\beta=0.25$ and $k=5$.
Figure 6. Fitted results for Principle 5 (small world). Top (a-c): Regression plots relating average shortest path length ($L$) and average clustering coefficient ($C$) with $n$ for $\beta\in\{0.25,0.5,0.75\}$ and $k=5$ for GPT-3.5. The value $a$ in legends represents the effect size (slope of the regression lines). Middle (d-f): Estimated values $\hat{\beta}$ of $\beta\in\{0.25,0.5,0.75\}$ for LLM-generated networks based on matching the average clustering coefficient and difference in the average shortest path between LLM-generated networks and Watts-Strogatz with the estimated rewiring probability $\hat{\beta}$ for GPT-3.5 agents. We report the $P$-values of the t-test comparing the average shortest path length of the LLM-generated networks and the average shortest path length of the Watts-Strogatz graphs with rewiring probability $\hat{\beta}$. Bottom (g): Regression plot for the relation $L\sim\log(n)$ for different LLM models and environments (school, work, community) for $\beta=0.25$ and $k=5$. The legend shows the effect size ($a$) and the $P$-value. (*: $P<0.05$; **: $P<0.01$, and ***: $P<0.001$.)